# Algorithmic Collusion And The Minimum Price Markov Game

### v1

@@ 5 Computer Experiments
We created a MARL environment following the MPMG formulation described in Section 4.2. In this setup, the MPMG is implemented as an episodic task, where each episode is one step long and corresponds to a single auction denoted as $t\in\{1,\dots,T\}$. Implementing the MPMG as an episodic task with one-step episodes aligns with Assumption 2 and simplifies the learning process by making state transitions and cumulative rewards straightforward. Implementation details are available in Appendix B. For the experiments, we set $\alpha=1.3$, meaning the collusive bid is 30% higher than the fair bid, as per the notations in Section 4.2.
When considering solution methods, techniques explicitly designed to foster cooperation, such as shared joint action spaces [47] and algorithmic communication [48], as well as those implicitly encouraging cooperation, like reputation-based models [49] and status-quo loss [50], are not suitable. Inducing cooperation in MARL settings is well-studied, but using these approaches would simulate scenarios where firms deliberately bias their pricing algorithms towards cooperation. Although this raises interesting questions, it is beyond the scope of this study. Nevertheless, we provide insights on this topic in Section 5.4. Here, we evaluate the inherent properties of the MPMG, requiring unbiased experiments achieved by assuming isolated firms. If collusion occurs in the MPMG, it would stem from the game’s inherent nature. Therefore, we used basic yet established state-of-the-art MARL algorithms. Additionally, since the artificial firms are considered isolated and do not benefit from sharing mechanisms (such as a common policy or state-value approximator), they remain symmetric in their algorithmic structures, supporting their decision-making processes. In other words, they share the same level of rationality.
In this section, we present the results of our computer experiments on the MPMG using three foundational MARL approaches: Multi-Armed Bandits (MAB), Deep Q-learning, and Actor-Critic Policy Gradient. As discussed in Section 2, MAB focuses on balancing the exploration-exploitation trade-off, Deep Q-learning aims at value function approximation for action selection, and Actor-Critic Policy Gradient integrates both value estimation and policy optimization. These methodologies represent the core families of techniques in MARL, providing a comprehensive perspective on their effectiveness in various scenarios.
A total of seven algorithms—three bandit variants, two Q-learning variants, and two policy gradient algorithms—were trained on four MPMG configurations: the 2-player homogeneous, the 5-player homogeneous , the 2-player heterogeneous, and the 5-player heterogeneous MPMG. In the heterogeneous cases, we set $\sigma(\beta)$ to 0.5. As discussed in Section 4, $0\leq\sigma(\beta)\leq 1$ represents the standard deviation of the power parameters ($\beta$) distribution, with $\sigma(\beta)=0$ representing the homogeneous case. Each algorithm was trained on 100 auctions over 100 replications for statistical robustness. Results are presented in graphs and tables in this section. Additional material, such as graphs about algorithmic convergence and other relevant metrics, can be found in Appendix A, while Appendix B contains implementation details (e.g., hyperparameter values).
5.1 Multi-Armed Bandits
Our first implementation involves representing the MPMG as a bandit problem because MAB algorithms are lightweight and can be used independently of a Markov game formulation. Thus, our first implementation with bandits resembles an iterated MPG, as we do not utilize the Markov property. Implementing MAB as a first step before tackling model-based MARL methods is advantageous due to its focus on the exploration-exploitation trade-off. MAB provides a low-cost approach to eventually uncovering core strategies, serving as a proof-of-concept essential to evaluating the MPMG. In this study, we used three bandit algorithms: $\epsilon$-greedy, Thompson Sampling (TS), and Upper Confidence Bound (UCB).
The $\epsilon$-greedy bandit algorithm is a simple yet effective strategy for balancing exploration and exploitation in multi-armed bandit problems. At each step, the algorithm selects a random action with probability $\epsilon$ (exploration) and the action with the highest estimated reward with probability $1-\epsilon$ (exploitation). This approach ensures that the algorithm occasionally explores new actions to discover potentially better rewards while primarily exploiting the best-known actions to maximize cumulative rewards.
TS is a Bayesian approach to the multi-armed bandit problem that balances exploration and exploitation by maintaining a probability distribution for the expected reward of each arm. At each step, it samples from these distributions and selects the arm with the highest sampled value. Over time, this method efficiently narrows down the best-performing arm while still exploring other options.
The UCB algorithm addresses the exploration-exploitation trade-off by selecting the arm with the highest upper confidence bound of the estimated reward. It adds a confidence interval to the estimated reward of each arm that decreases as more data is collected, encouraging the exploration of less-tried arms initially and favoring exploitation as more information is gathered.
The performance of our bandit algorithms is displayed in Figure 5.1 in terms of training joint action frequencies, which represent the strategy profiles played during training. Since bandit agents are stateless, presenting the results of evaluation episodes is unnecessary, as their strategies would not evolve as the game unfolds. By evaluation, we mean making agents play with their learned and crystallized policies. Figure 5.1 (a) and (b) demonstrate that both $\epsilon$-greedy and TS agents have learned to play the Nash equilibrium strategy, which involves bidding the fair price. This outcome is expected, as bandit algorithms are designed to optimize immediate rewards and naturally gravitate towards strategies that consistently yield the highest payoff. In the context of the iterated MPG, defecting is the dominant strategy, leading to higher individual payoffs when both agents defect. However, Figure 5.1 (c) shows that UCB agents consistently managed to coordinate towards collusion, with their policies translating into a 60% frequency for $(CP,CP)$ on average. As expected, the frequencies of $(CP,FP)$ and $(FP,CP)$ decrease as agents exploit more, reaching marginal values, while the Nash equilibrium strategy ranks second with a 35% frequency on average after 100 auctions.
Figure 5.1: Training joint action frequencies of $\epsilon$-greedy (a), TS (b), and UCB (c) on the 2-player iterated homogeneous $(n=2,\sigma(\beta)=0)$ MPG over 100 replications of 100 iterations each. The colored bands show the confidence intervals over 100 replications.
(a) $\epsilon$-greedy
(b) Thompson Sampling
(c) UCB
Moreover, as expected, $\epsilon$-greedy and TS agents did not manage to learn how to collude in the $n$-player heterogeneous MPMG (see Appendix A), as the number of agents $n$ and the level of heterogeneity $\sigma(\beta)$ are known to hinder the collusive potential of a market as coordination gets harder. Consequently, we also expect UCB agents to learn to play the Nash more often as $n$ or $\sigma(\beta)$ increases. In fact, this is what we observe in Figure 5.2. We can see that market heterogeneity and the number of agents both negatively and independently impact the market’s collusive potential. Note that Figure 5.2 (b) does not display any legend for clarity since there are $2^{n}$ possible joint actions; however, the red and grey plots, respectively, represent the Pareto dominant strategy and the Nash, being consistent with Figure 5.2 (a). As we can see from Figure 5.2 (c), not only market heterogeneity and the number of agents force UCB agents to shift towards the Nash, but when both factors are set to inconvenient values, the uncertainty increases.
Figure 5.2: Training joint action frequencies drawn by UCB agents in the 2-player heterogeneous (a), the 5-player homogeneous (b), and the 5-player heterogeneous iterated MPG (c). The colored bands show the confidence intervals over 100 replications.
(a) $n=2$, $\sigma(\beta)=0.5$
(b) $n=5$, $\sigma(\beta)=0$
(c) $n=5$, $\sigma(\beta)=0.5$
Since a residual amount of collusion can occur during training regardless of the agents’ direction, we have set a threshold to determine what can be called collusion. This threshold is 30% of the maximal collusive spoil possible. TS and $\epsilon$-greedy agents have learned the Nash equilibrium 100% of the time, and they never achieved a significant amount of collusion. However, UCB agents, in regard to our threshold, managed to collude 100% of the time on the 2-player homogeneous iterated MPG. Table 1 reports the statistics related to collusion for UCB agents, including the collusion rate, which tells how many times UCB agents have achieved a collusive spoil that exceeds the threshold.
Table 1: UCB agents’ performance regarding collusion on 100 training auctions over 100 replications. The collusion rate is given as a percentage of the number of replications where collusion exceeded the threshold of 30% of the potential maximal collusive spoil. The average, standard deviation (STD), max, and min spoil are given as a percentage of the maximal collusive spoil achievable. These statistics are given for several market configurations.
$n=2$, $\sigma(\beta)=0$ $n=2$, $\sigma(\beta)=0.5$ $n=5$, $\sigma(\beta)=0$ $n=5$, $\sigma(\beta)=0.5$
Collusion Rate 100 4 40 0
Average Spoil 64.5 9 25 0
STD Spoil 1.5 8.8 11.6 2.6
Max Spoil 66 59 57 15
Min Spoil 63 0 17 0
Table 1 shows that market heterogeneity impedes collusive potential more than the number of agents, as UCB agents have achieved collusion 40% of the time in the 5-player homogeneous MPG and only 4% of the time in the 2-player heterogeneous MPG. Moreover, the average collusive spoil is much higher in the 5-player homogeneous case than in the 2-player heterogeneous case. When both market heterogeneity and the number of agents increase, collusion is fully tamed, as the average collusive spoil in this case is 0% of the maximal spoil achievable.
5.2 Deep Q-learning
Assuming isolated firms, each agent can treat competitors as part of the environment and learn a Q-value function independently, considering the impact of other agents’ actions as part of the state transitions. This leads to the following representation for a given state:
$Q_{i}(s,a_{i})=\mathbb{E}\left[R_{i}\mid s,a_{i}\right],$ (5)
where $s$ encodes information about all players. When opponent modeling is used, opponents’ behaviors are explicitly incorporated into the Q-value function. The enhanced Q-value function, denoted as $Q(s,a,s_{o},a_{o})$, incorporates the opponent’s state $s_{o}$ and action $a_{o}$ directly into the learning process. For example, with 2 agents
$Q(s,a,s_{o},a_{o})=\mathbb{E}[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime},s^{\prime}_{o},a^{\prime}_{o})|s,a,s_{o},a_{o}],$ (6)
where $s$ and $s_{o}$ are the states of the agent and the opponent, respectively, $a$ and $a_{o}$ are their corresponding actions, $r$ is the reward, $\gamma$ is the discount factor, and $a^{\prime}$ and $a^{\prime}_{o}$ are the next actions of the agent and the opponent.
In our experiments, we used the Double Deep Q-Network (D3QN) [51] in both the versions depicted by Equation (5) and (6). We refer to the former case as naive D3QN, and the latter as D3QN with opponent modeling (D3QN-OM). Figure 5.3 shows that the MPMG is robust to tacit coordination among rational agents. The naive D3QN may be naive in its structure compared to the D3QN-OM, but it still encodes a higher level of rationality than bandit algorithms that do not use state representation. From Figure 5.3 (a), we can see that D3QN agents converge with good certainty towards the Nash equilibrium. Indeed, D3QN agents learn to assign higher Q-values to action 0 as the game unfolds, as described by Figure 5.3 (b) and (c). It follows without surprise that the naive D3QN agents have learned to play along the same pattern when either the number of agents or market heterogeneity increases (see Appendix A).
Figure 5.3: Joint action frequencies of the naive D3QN during evaluation (a), training Q-values for action 0 (b), and training Q-values for action 1 (c) in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation episodes
(b) Q-values for action Fair Price during training
(c) Q-values for action Collusive Price during training
It turns out that the impact of opponent modeling on Q-agents’ ability to discover profit-maximizing strategies in the 2-player homogeneous MPMG is small. Indeed, D3QN-OM agents achieve some amount of collusive spoil, which is more than the naive D3QN agents did. However, as we can observe from Figure 5.4 (a), D3QN-OM agents seem uncertain. Indeed, the Nash equilibrium has an average rate of near $0.5$, and its associated confidence interval is rather wide. Nevertheless, D3QN-OM agents, on average, converge towards the Nash equilibrium, as shown in Figure 5.4 (b) and (c). Additionally, as with the naive D3QN, we observe no surprises regarding the agent behavior in the general $n$-player heterogeneous setting (see Appendix A).
Figure 5.4: Joint action frequencies of the D3QN-OM agents during evaluation (a), training Q-values for action 0 (b), and training Q-values for action 1 (c) in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation episodes
(b) Q-values for action Fair Price during training
(c) Q-values for action Collusive Price during training
Despite the fact that D3QN agents never passed the collusive threshold, their collusive potential exists, and the D3QN-OM method is somewhat an improvement from the naive D3QN as we can see in Table 2.
Table 2: D3QN-OM agents’ performance regarding collusion on 100 training auctions over 100 replications. The collusion rate is given as a percentage of the number of replications where collusion exceeded the collusive threshold of 30% of the potential maximal collusive spoil. The average, standard deviation (STD), max, and min spoil are given as a percentage of the maximal collusive spoil achievable. These statistics are given for several market configurations.
$n=2$, $\sigma(\beta)=0$ $n=2$, $\sigma(\beta)=0.5$ $n=5$, $\sigma(\beta)=0$ $n=5$, $\sigma(\beta)=0.5$
Collusion Rate 0 0 0 0
Average Spoil 17.6 17.5 0 0
STD Spoil 11.5 11.1 0 0
Max Spoil 30 33 8 0
Min Spoil 0 0 0 0
5.3 Actor-Critic Policy Gradient
In the last experiments, we used Multi-Agent Proximal Policy Optimization (MAPPO), a policy gradient algorithm where the policy represents the probability distribution over bidding either the fair price or a collusive price. MAPPO employs the actor-critic method, in which a critic evaluates the actions taken by the actor by estimating the value function—representing expected returns from the current state following the current policy. The actors are updated based on the gradients of the expected reward, as assessed by the critic. The goal of the actors is to refine the policy using feedback on their performance from the critic.
Figure 5.5: Joint action frequencies during evaluation (a), training policy for action 0 (b), and training policy for action 1 (c) of MAPPO agents in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation
(b) Policy for action Fair Price during training
(c) Policy for action Collusive Price during training
We found that MAPPO agents always learn to play the Nash, regardless of market settings. As suggested by Figure 6.3, which shows the MAPPO results on the 2-player homogeneous MPMG, tacit coordination does not occur in the $n$-player heterogeneous setting (see Appendix A). Figure 6.3 (b) and (c) show that MAPPO agents neatly and consistently converge towards the Nash play.
@@ 5.4 Conclusion
The results of our experiments underscore the varying effectiveness of different MARL approaches in the MPMG. From our observations in Section 5, three key insights emerge.
First, there appears to be no positive correlation between the level of complexity of our AI agents and the emergence of tacit coordination. Interestingly, only the UCB agents, utilizing a relatively simple method, consistently achieved substantial collusive spoils while demonstrating robustness in non-favorable market conditions (see Table 1). This may seem counterintuitive, as greater algorithmic complexity and knowledge are typically expected to provide more perspective. However, it is important to note that the rational strategy is to play the Nash equilibrium, as humans tend to do in similar settings [52]. Therefore, we can draw the following partial conclusion: assuming the MPMG reasonably approximates minimum price-ruled auctions, the minimum price rule is resilient to non-engineered tacit coordination among rational actors, though paradoxically, collusion can occur in the most naive settings.
Second, the emergence of tacit coordination does not appear to rely heavily on self-reinforcing trends. The state representation, evolving linearly as the game progresses, seems to favor the Nash equilibrium. This can be attributed to the fact that achieving the profit-maximizing Pareto equilibrium requires a fragile and precise set of conditions if it depends on self-reinforcing trends. In practice, players will always respond to cues of defection from their opponents, and since agents are symmetrical (at least in this study), no player is likely to make the first move. Coordination could theoretically be triggered by sustained mutual exploration of defection, which is even less probable when agents are fully isolated, as is the case here. The question of whether tacit coordination arises from self-reinforcing empirically-based beliefs or from planned behavior supported by encouraging beliefs about opponents remains open.
Third, the MPMG exhibits properties consistent with Jean Tirole’s seminal work on market collusion potential [53]. Specifically, as the number of firms increases, coordination becomes more challenging because the value of $u_{i}(CP,k=0)$ decreases, undermining the intrinsic value of the Pareto dominant strategy and thus reducing the likelihood of successful collusion. Additionally, market heterogeneity, characterized by varying market power or market share among firms, negatively impacts the incentive to coordinate. This supports the idea that the MPMG provides a simple yet realistic approximation of market dynamics.

### v2

@@ 4 Computer Experiments
Cooperation in multi-agent reinforcement learning (MARL) settings is well studied. However, this algorithmic cooperation mostly relies on explicit designs—such as shared joint action spaces [48], algorithmic communication [49], reputation-based models [50]—which are not suitable for our context, as these approaches emulate scenarios where agents deliberately bias their actions towards cooperation.
In this study, we evaluate the inherent properties of the MPMG, requiring unbiased players. If collusion occurs, it would arise solely from the game’s intrinsic structure. Therefore, we employed basic yet established state-of-the-art MARL algorithms to define artificial agents that are symmetric in their decision-making processes, do not communicate, and do not hold beliefs about their opponents. In other words, we aim to create isolated and equally rational agents.
We conducted a total of twenty-four experiments, combining six agent types with four MPMG configurations (2-player homogeneous, 5-player homogeneous, 2-player heterogeneous, and 5-player heterogeneous MPMG). As discussed in Section 3, $0\leq\sigma(\beta)\leq 1$ represents the standard deviation of the power parameters ($\beta$) distribution, with $\sigma(\beta)=0$ representing the homogeneous case. For experiments in the heterogeneous setting, we set $\sigma(\beta)=0.5$. Additionally, we set $\alpha=1.3$, meaning that the collusive bids were 30% higher than the fair bids. Each agent was trained on 100 auctions over 100 replications for statistical robustness.
4.1 Agent Representation
We based our agents on three foundational approaches of MARL, that is, Multi-Armed Bandits (MAB), Deep Q-learning, and Actor-Critic Policy Gradient. Implementation details are available in Appendix B.
Definition 2.
Multiagent Reinforcement Learning (MARL). MARL extends single-agent reinforcement learning (RL) techniques to settings with multiple agents learning concurrently in a shared environment. In RL, an agent interacts with its environment to maximize cumulative rewards, where key components include agent, environment, state, action, reward, policy, and value function. The agent, acting as the decision-maker, responds to the environment, which consists of external factors and situations represented as states. Actions are possible decisions the agent can take, evaluated by rewards as feedback on their success. The policy is the agent’s strategy for selecting actions based on the current state, while the value function estimates future rewards to determine the desirability of states or actions. In MARL, multiple agents simultaneously learn and update their policies, causing the environment to appear non-stationary from any single agent’s perspective. This dynamic leads to more complex interactions compared to single-agent RL, requiring robust techniques to address the challenges of Markov games and multiagent systems in general.
The MPG as a bandit problem.
MAB algorithms [51] represent a simplified form of RL or MARL, focusing on the exploration-exploitation trade-off without considering future states (stateless), that is, without a model of the environment. MAB methods solve the bandit problem, where the decision-maker must sequentially choose among different actions with unknown reward distributions in order to maximize the total reward over time. Each action represents a distinct stochastic process with its own, a priori unknown, probability distribution of rewards. The goal is to minimize regret, which is the difference between the total reward the decision-maker could have earned by always choosing the best action and the actual reward obtained by balancing exploration and exploitation. MAB agents actually play an iterated MPG, as they do not utilize the observation space. We deployed $\epsilon$-greedy agents [52] that balance exploitation and exploration by selecting a random action with a probability $\epsilon$ decreasing over time, Bayesian Thompson Sampling (TS) agents [53] that maintain a probability distribution for the expected reward of each action, and Upper Confidence Bound (UCB) agents [54] that select the action with the highest upper confidence bound of the estimated return.
Deep Q-learning.
Methods based on Q-learning [55] maintain a state-action pair values, named Q-values, based on the state-action pair expected reward. Decision making is then conducted following the highest Q-values associated to state-action pair in question. In our experiments, we used the Double Dueling Deep Q-network (D3QN) algorithm [56]. In the D3QN, a multi-layer perceptron is used to approximate the Q-values, and the Double Q-learning approach is employed to reduce overestimation bias by maintaining two sets of Q-value estimates. The Dueling architecture, on the other hand, decomposes the Q-value into two separate streams representing the state-value function and the advantage function, which helps the model learn better representations of the value of being in a given state, independent of the specific actions.
Actor-critic policy gradient.
Unlike Q-learning, policy gradient methods focus on optimizing a parameterized policy $\pi_{\theta}(a|s)$ directly by gradient ascent on expected rewards [55]. In the actor-critic framework, two function approximators are employed. The actor, which is responsible for selecting actions based on the current policy, and the critic, which evaluates the action-value function to provide feedback on the quality of the chosen actions. The actor updates the policy in the direction suggested by the critic, while the critic learns to estimate the expected return, enabling a stable and efficient training process by reducing the variance of the policy gradient estimates. Here, we used Multi-Agent Proximal Policy Optimization (MAPPO) agents [57], that extend the Proximal Policy Optimization (PPO), a popular policy gradient method, to multi-agent settings. The core idea of PPO is to strike a balance between policy improvement and preventing drastic updates to the policy, which could harm performance. To achieve this, PPO uses a modified objective function that constrains how much the new policy is allowed to diverge from the old policy during an update.
4.2 Results
Before presenting the results of the experiments, it is essential to outline several important considerations for evaluating artificial agents in the multi-agent context. Indeed, the interpretation of the results depends on both individual and collective convergence.
By individual convergence, we mean the convergence of the algorithms themselves. For D3QN and MAPPO, this is assessed by examining training loss function values over time. In the case of MAPPO, both the actor and critic networks must converge. In our experiment, they all converged, as can be seen in Figure 4. Additionally, D3QN and MAPPO offer internal metrics indicating their policies. For MAPPO, we can directly observe the evolution of the policies regarding actions FP and CP, while D3QN uses Q-values for each action. The evolution of MAPPO policy values and D3QN Q-values during training is shown in Figure 5. In the case of the stateless bandit algorithms ($\epsilon$-greedy, TS, and UCB), cumulative regrets are generally used to establish learning in the single-agent context, though they are less meaningful in Markov games. Nevertheless, cumulative regrets are presented in Figure 6. Figures 4, 5, and 6 are provided in Appendix A.
However, the joint action frequencies, or the frequencies at which each strategy profile is played, are metrics common to all experiments, and they provide a clear picture of collective convergence. Therefore, they offer a reliable way to evaluate and compare our MARL agents in terms of behavioral trends. In the single-stage MPMG (MPG) or in similar static games, three distinguishable collective outcomes emerge: agents play the Nash profile, the Pareto profile, or they diverge in their choices, meaning that only a subgroup of players share a nonzero return associated with the Nash play. In dynamic settings, outcomes are based on convergence toward a particular outcome, with agents represented by stochastic processes typically converging toward mixed strategies. The challenge, then, is fitting a discrete interpretation (do they cooperate or not?) onto a continuous spectrum of behaviors.
The evolution of the average joint action frequencies over training episodes is displayed in Figure 3 in Appendix A. The average joint action frequencies from the last training episode for all experiments and outcomes are displayed in Table 2, showcasing the impact of atomicity and heterogeneity on the agents’ ability to coordinate, as well as the different behavioral trends emerging from each configuration of agents and market parameters.
Table 2: Average joint action frequencies across repeats at the last training episode for each outcome in each experiment. Bold values indicate the highest average joint action frequency among the three outcomes.
$n=2$, $\sigma(\beta)=0.0$ $n=2$, $\sigma(\beta)=0.5$ $n=5$, $\sigma(\beta)=0.0$ $n=5$, $\sigma(\beta)=0.5$
FP CP Other FP CP Other FP CP Other FP CP Other
D3QN 0.37 0.16 0.47 0.37 0.17 0.47 0.07 0.02 0.91 0.07 0.02 0.91
TS 0.72 0.04 0.24 0.69 0.04 0.27 0.41 0.01 0.59 0.32 0.01 0.67
$\epsilon$-greedy 0.70 0.03 0.27 0.70 0.03 0.27 0.35 0.01 0.64 0.35 0.01 0.64
MAPPO 0.52 0.10 0.38 0.51 0.11 0.39 0.24 0.01 0.75 0.21 0.01 0.78
UCB 0.40 0.58 0.02 0.43 0.55 0.02 0.34 0.43 0.23 0.22 0.08 0.70
In addition, Figure 1 ranks each agent-MPMG configuration pair in terms of collusive potential by using scores based on the data in Table 2. More specifically, each value in Figure 1 represents the difference between the average joint action frequency of the last training episode of the Nash and Pareto outcomes. The resulting value is then scaled from 00 (most Nash) to $1$ (most Pareto) using the minmax formula.
Figure 1: One-dimensional heat-map highlighting the relative difference among agents and MPMG configurations in terms of their collusive potential. From the most Nash (left) to the most Pareto (right) configuration.
@@ 4.3 Conclusion
Our results, summarized in Table 2 and Figure 1, underscore the varying effectiveness of different MARL approaches in the MPMG. We can see that, in the $5$-player heterogeneous MPMG, and in contrast to the results of the $2$-player homogeneous MPMG, agents always converge towards a suboptimal outcome far from the Pareto profile, but closer to the Nash Equilibrium. This shows the impact of atomicity and heterogeneity on the capacity of agents to coordinate themselves, let alone cooperate in collusion.
Also, UCB agents seem to favor the Pareto profile, consistently across different configurations. In fact, UCB is the only algorithm that not only fosters higher rates of convergence towards the Pareto profile but also shows robustness against atomicity and heterogeneity, at least when they are increased individually. This is clear in Figure 1, where we can see that the three highest scores belong to UCB agents. However, we can see in Table 2 that in the $5$-player heterogeneous MPMG, UCB agents behave similarly to their counterparts.
Moreover, it seems there is no correlation between algorithmic sophistication and collusive potential. Stateless bandits can either converge towards an approximate Nash Equilibrium (TS, $\epsilon$-greedy) or favor the Pareto outcome (UCB), while methods using more information either favor the Nash Equilibrium (MAPPO) or other suboptimal outcomes (D3QN).